OpenStudy (anonymous):

integral from 0 to t... k, M, v, sin(theta), g are all constants

OpenStudy (anonymous):

\[\int\limits_{0}^{t}[e^{-kt/M}(-{k \over M}v \space \sin \theta - g)+g]dt\]

OpenStudy (anonymous):

for simplification... let\[-{k \over M}v \space \sin \theta= \tau \]

OpenStudy (anonymous):

\[\int\limits\limits_{0}^{t}[e^{-kt/M}(\tau - g)+g]dt\]

OpenStudy (anonymous):

\[\int\limits\limits\limits_{0}^{t}[e^{-kt/M}\tau - e^{-kt/M}g+g]dt\]

OpenStudy (anonymous):

no longer need help